# Judging the Judges: A Systematic Study of Position Bias in LLM-as-a-Judge

### v7

@@ 4 Main Findings
Position Bias of Capable Judges are not Mere Random Variations: As shown in Table 2, the capable judges on the benchmark tasks, supported by minimal "Error" rates, generally exhibit $RS$ values above 0.85. The most capable models, such as Claude-3.5-Sonnet, Claude-3-Opus, GPT-4, and GPT-4o, all achieve near-perfect $RS$ scores exceeding 0.95 on both benchmarks. These results confirm that LLM judgments, and the resulting position bias, are not due to random variations. This strengthens confidence that one-time generated judgments by these validated LLMs accurately reflect their judging capabilities.
Position Bias Varies by Judge & Task: As shown in Table 2, in general, GPT-4 excels as the best-performing judge model across metrics and benchmarks, hence considered the baseline model in Fig. 2 (a) for $PC$ comparisons. Moreover, certain models achieve comparable or superior performances than GPT-4 on certain tasks. For example, according to Fig. 2 (a), for coding task evaluation, GPT-4o and Claude-3.5-Sonnet are likely more ideal judges. Also, GPT-3.5-Turbo achieves comparable $PC$ as GPT-4, indicating that it could be employed as a cost-effective alternative to coding evaluations.
Additionally, significant variations of $PF$ across tasks are observed. The judges achieving close-to-0 $PF$ in general, such as GPT-4 and Claude-3.5-Sonnet, exhibit varied preference directions across tasks, preferring primacy on some tasks while recency on others. Particularly, o1-mini, while being primacy-preferred on coding, extraction, and math, exhibits almost fair preferences on reasoning, role play, and writing tasks. Even for judges that are recency-preferred across all tasks (e.g., Claude-3’s and Gemini-pro’s), the extent of biased preference, as reflected by $PF$ values, varies by task.
Moreover, a high position consistency does not guarantee fairness. For example, on coding task evaluations, GPT-4 and GPT-4o achieve the top consistency but are significantly recency-preferred and primacy-preferred, respectively. In comparison, GPT-3.5-Turbo is highly preference fair while having comparable consistency.
@@
In list-wise comparisons, similar variations by judge and task in position bias were observed, as shown in Fig. 2 (b). Furthermore, Table 2 highlights that more capable models, such as GPT-4o and Claude-3.5-Sonnet, maintain high consistency when transitioning from pairwise to list-wise evaluations, while less capable models, such as GPT-3.5-Turbo, exhibit greater sensitivity to the increased number of candidates in list-wise tasks.
@@
Therefore, the position bias of LLM judges is judge-dependent and task-dependent. This observation is further confirmed by the bidirectional stepwise regression where judge identities and task categories are statistically significant predictors of $PC$ and $PF$. In practice, to achieve the most reliable evaluations, one needs to select task-specific LLMs whose consistency and fairness are in balance. Similar findings can be observed in the DevBench results, as detailed in Appendix.Sec. E.2.
@@
Position Bias Correlates to Answer Quality Gap: Intuitively, the difficulty of judging a pair of candidate answers is largely reflected by their difference in quality. In this study, as defined in Section 2.3, we quantify the quality gap ($\delta_{q}$) between candidate solutions and expected baseline (calculated by $1/p$ for a $p$-candidate list) by the overall win rate ($owr$). Therefore, $\delta_{q}$ increases as $owr$ extends from baseline to 0 or 1. Fig. 2 (e) and (f) exhibit significant parabolic shapes, indicating that $PC$ is positively proportional to $\delta_{q}$. This aligns with our intuition that the answers with larger quality disparities are easier to achieve judgment consistency, whereas those of similar quality are difficult to judge, increasing the likelihood of position bias that leads to lower $PC$. The same relationship is observed for each individual judge and across benchmarks, as demonstrated in Appendix.Sec. E.
@@
Together, we conclude that as the answer quality gap enlarges, judges generally become more position consistent and preference fair according to the regression curves. However, exceptions are common, as shown by the individual scatter points of these figures. This indicates that though the answer quality gap significantly influences the position bias of LLMs, other factors also play important roles. Therefore, built on our findings, future studies may have better control over the answer quality gap when evaluating LLM judges, exploring other impacting factors on position bias, and seeking potential mitigation strategies.
@@
LLM Agreement Analysis: We complement our investigation of position bias with an agreement and disagreement analysis among LLM judges. Rather than focusing exclusively on overall consistency or fairness, we examine how judges converge and diverge in their assessments of individual instances. Agreement analysis quantifies the percentage of instances where two judges mutually agree on the outcome. Disagreement analysis counts the number of choices deviating from the mode for each instance among all judges. This further complies a “distribution of disagreement” across the dataset.
Our findings reveal that, despite exhibiting similar overall $PC$ and $PF$ scores, judges vary significantly in their judgments on individual instance. Disagreement analysis, in particular, highlights instances where consensus is either easily or difficultly achieved, reflecting the inherent complexity of the judgment task. For example, as shown in Fig. 3, more than half of the dataset can be considered relatively easy to judge, as over 75% of all 12 judges agree with each other on these instances. Conversely, fewer than 4% of instances represent the likely hard-to-judge cases where a majority of judges fail to reach consensus.
Figure 3: Distribution of disagreement on MTBench. The y-axis indicates the proportion of the dataset where the level of disagreement among judges does not exceed a specific threshold.
Based on our observations of answer quality gaps and LLM agreement/disagreement patterns, this study offers practical insights for designing evaluator benchmarks that account for the varying difficulty levels of judgment tasks. Specifically, the most challenging instances to evaluate are characterized by: (1) frequent disagreements among LLM judges, (2) closely matched win rates and minimal quality gaps among candidate models, and (3) significant position bias exhibited by the majority of judges. Further discussions and analyses can be found in Appendix Sec. C.
@@ 6 Limitations
Despite proposing scalable metrics and investigating key factors influencing position bias, our study has several limitations. First, due to computational constraints, we evaluated only twelve commercial LLM judges for pairwise settings and four for list-wise settings, limited to 3-candidate lists, using two benchmarks. Open-source models, such as Llama 7B-13B, struggled to provide reliable judgments, restricting their inclusion. Additionally, while we tested original benchmark prompt templates, exploring variations in prompting techniques could provide further insights. Future work could expand this study to include more models, tasks, prompting strategies, and larger list-wise candidate pools to validate and generalize our findings.
Second, data accessibility limitations prevented a direct analysis of Judge-level factors like architecture and parameter size. Instead, we approximated these factors by grouping models by family properties. Open-source models with accessible architectural details could allow deeper analysis of these influences. Additionally, our analyses rely on completed judgments, such as measuring the answer quality gap and task difficulty after evaluations. Future work could explore methods to estimate or control these factors pre-judgment, reducing computational costs and enabling proactive mitigation strategies.
Lastly, our focus was on understanding position bias rather than mitigating it. While our findings provide a foundation for effective mitigation, further research is needed to address issues like maintaining consistency and fairness when answer quality gaps are minimal, where position bias is most pronounced. Multivariate analyses exploring interactions between factors like prompt length, task complexity, and answer quality gaps could also yield deeper insights and enhance mitigation approaches.
@@ Appendix A Reproducibility
Our experiments were conducted primarily using API access, with a total cost of approximately 3,000 USD.
The code repository for reproducing our results is available at: https://anonymous.4open.science/r/Position-Bias-Analyzer-Demo-F7E3
@@ Appendix C LLM Agreement Analysis
Figure 4: Disagreement Analysis and Mutual Agreement Heatmap of LLM judges on MTBench and DevBench. The blue bars of Figures (a) and (b) represent the number of instances with certain disagreement values, and the red cumulative curve displays the cumulative count and corresponding percentage over the benchmark dataset. The heat maps of Figures (c)(d)(e) marked higher mutual agreement values with a brighter color.
@@ C.1 Mutual Agreement & Familial Property
We compute the LLM judges’ mutual agreement on the instances to explore how “alike” or consistent they are across a set of evaluations. We denote two judges agree on an instance if their judgment choices are identical. Then the mutual agreement between two judges on a benchmark is defined as the proportion of their agreed instances. Fig. 4(c)(e) displays the mutual agreement heatmap for all judges on MTBench and DevBench, respectively. For MTBench that utilizes the 3-option mode, we also consider the “without tie" agreement since two judges are less disagreed when one chooses {C} while the other prefers a certain solution, compared to the case when they prefer different solutions. The “without tie” agreement heatmap of the twelve judges on MTBench is explored in Fig. 4(d).
@@ C.2 Disagreement & Benchmark Design Insight
Since the mutual agreement between LLM judges is not perfect and usually a considerable proportion of instances are difficult for them to reach a consensus, disagreement analysis becomes crucial and insightful. Therefore, we define the disagreement of an evaluation instance to be the number of judgments different from the majority. By this definition, an instance with all judges reaching a consensus on the better solution will have a disagreement of 0; in contrast, an instance where judgments are widely varied will result in a high disagreement. For our study where twelve judges are investigated, the maximum disagreement of an MTBench instance is 8, accounting for the 4{A}-4{B}-4{C} choice pattern by 3-option mode. On the other hand, for DevBench instances, the maximum possible disagreement is 6, representing the 6{A}-6{B} judgment distribution for the 2-option mode.
The distributions of instances with different disagreement values on MTBench and DevBench are shown in Fig. 4(a) and Fig. 4(b), respectively. From our disagreement analysis, at least 75% of the judges reached a choice consensus on more than half of the instances on both benchmarks. These are likely easy-to-evaluate instances, and the reliability of LLM judgments is enhanced by majority voting. In comparison, the instances with the highest disagreement are likely the ones that are difficult to evaluate and where the position bias is most likely to occur. However, luckily, these instances are rare, occupying only 0.03% and 0.07% of MTBench and DevBench respectively. In other words, majority voting of multiple capable LLM judges could be practically useful for over 99% of evaluation instances on both benchmarks.
Moreover, if we roughly consider the disagreement value of instances as their difficulty for judging, then Fig. 4(a) and Fig. 4(b) exhibit a balanced distribution of instances with varied difficulty. This is because, except for the instances with the highest disagreement, the numbers of other instances with varied disagreement do not vary significantly, indicating a smoothly increasing difficulty curve across the benchmark datasets.
@@
Appendix D Open-source Model Discussion
Due to computational constraints, we conducted limited experiments on open-source models, specifically Llama 7B-13B, as representatives of smaller-scale models. However, as shown in Table 2, these models struggled to perform effectively on our evaluation tasks. A primary issue was their limited context window, which proved insufficient for benchmarks like DevBench, where questions often include extensive instructions, such as full product requirement documents. This context limitation caused frequent truncation of either the prompt or the candidate responses, leading to incomplete or erroneous outputs and significantly higher error rates.
Figure 5: Llama-13B output example: Despite providing analysis and a choice, the output format does not match the specified “[[choice]]" pattern, causing extraction failures.
Another key factor affecting performance was mismatched output formatting. MTBench requires responses in the format [[choice]] (e.g., [[A]]), but Llama-13B, as shown in Fig. 5, frequently produced outputs such as [A], which could not be parsed correctly by the regular expression used for choice extraction. Interestingly, Llama-7B outperformed Llama-13B on MTBench, likely due to its simpler, shorter responses that inadvertently aligned better with the expected format. These errors illustrate the challenges these models faced in meeting benchmark-specific requirements, highlighting nuances in their evaluation results. Considering practicality and effectiveness, we only focus on more capable commercial judge models.
Appendix E More Results of Position Bias and Answer Quality Gap Measurement
E.1 MTBench
@@ E.2 DevBench
Figure 7: Baseline comparisons of judges across tasks on DevBench. An asterisk marks the statistical significance by Student’s t-tests. Figure (a) includes a radar chart comparison by family and an overall baseline bar chart comparison of $PC$. The baseline model is chosen to be GPT-4 for its generally superior performances. Figure (b) demonstrates the $PF$ score of evaluated judges across tasks where the baseline is $PF=0$.
Appendix F Variable Selection and Tests
F.1 Bidirectional Stepwise Regression with AIC
@@ F.2 Test results
Appendix G Experiment Settings
@@
G.1 Judges, Candidates, and Tasks

### v8

@@ 4 Main Findings
Figure 2: Judge performances on MTBench. Fig. (a)(b) are the radar charts for the $PC$ comparison by family, judge, and task. Fig. (c) leverages linear regression to explore the general relationship between $PC$ and $PF$. Fig. (d) to (f) investigate the impact of answer quality gap on position bias using overall win rates. Fig. (a) to (e) are for pairwise comparative settings, while Fig. (f) are obtained under list-wise evaluations.
@@
Position Bias of Capable Judges are not Mere Random Variations: As shown in Table 2, the capable judges on the benchmark tasks, supported by minimal "Error" rates, generally exhibit $RS$ values above 0.85. The most capable models, such as Claude-3.5-Sonnet, GPT-4, and Llama-3.3-70B, all achieve near-perfect $RS$ scores over 0.95 on both benchmarks. These results confirm that judgments from capable LLM judges, and the resulting position bias, are not due to random variations. This strengthens confidence that one-time generated judgments by these validated LLMs accurately reflect their judging capabilities.
Position Bias Varies by Judge & Task: As shown in Fig. 2(a), position bias among LLM judges varies significantly across different judges and tasks. For instance, GPT-4o demonstrates higher position consistency when evaluating coding tasks but performs less consistently on other tasks compared to GPT-4. Similarly, Gemini-1.5-pro, while achieving higher $PC$ than other Gemini models in most tasks, exhibits comparable consistency when judging extraction tasks. Similar findings can be observed in the DevBench results, as detailed in Appendix.Sec. D.2.
Variations in preference fairness are also evident. As shown in Table 2, GPT-4 and GPT-3.5-Turbo display different preference biases across datasets and tasks—being recency-preferred on MTBench but primacy-preferred on DevBench. Likewise, Claude-3.5-Sonnet, which is nearly preference-fair on MTBench ($PF=0.01$), exhibits a strong recency-preferred position bias on DevBench ($PF=0.22$).
While higher position consistency generally correlates with improved preference fairness (as demonstrated by the regression curve in Fig. 2(c)), consistency alone does not guarantee fairness. Certain LLM judges, despite achieving high $PC$, still exhibit significant and varied preference directions across different tasks, underscoring the need to evaluate both consistency and fairness when assessing LLM judges.
@@
In list-wise comparisons, similar variations in position bias were observed across judges and tasks. Furthermore, Table 2 highlights that more capable models, such as GPT-4o and Claude-3.5-Sonnet, maintain high consistency when transitioning from pairwise to list-wise evaluations, while less capable models, such as GPT-3.5-Turbo, exhibit greater sensitivity to the increased number of candidates in list-wise tasks.
@@
Therefore, the position bias of LLM judges is both judge-dependent and task-dependent. This observation is further confirmed by the bidirectional stepwise regression where judge identities and task categories are statistically significant predictors of $PC$ and $PF$. In practice, when evaluating LLM judge’s reliability or selecting suitable LLM judges, considering the balance between consistency and fairness, as well as accounting for task-specific variations, may be beneficial.
@@
Position Bias Correlates to Answer Quality Gap: Intuitively, the difficulty of judging a pair of candidate answers is largely reflected by their difference in quality. In this study, as defined in Section 2.3, we quantify the quality gap ($\delta_{q}$) between candidate solutions and expected baseline (calculated by $1/p$ for a $p$-candidate list) by the overall win rate ($owr$). Therefore, $\delta_{q}$ increases as $owr$ extends from baseline to 0 or 1. Fig. 2 (e) and (f) exhibit significant parabolic shapes, indicating that $PC$ is positively proportional to $\delta_{q}$. This aligns with our intuition that the answer pairs or series with larger quality disparities are easier to achieve judgment consistency, whereas those of similar quality are difficult to judge, increasing the likelihood of position bias that leads to lower $PC$. The same relationship is observed for each individual judge and across benchmarks, as demonstrated in Appendix.Sec. D.
@@
Together, we conclude that as the answer quality gap enlarges, judges generally become more position consistent and preference fair according to the regression curves. However, exceptions are common, as shown by the individual scatter points of these figures. This indicates that though the answer quality gap significantly influences the position bias of LLM judges, other factors also play important roles. Therefore, built on our findings, future studies may have better control over the answer quality gap when evaluating LLM judges, exploring other impacting factors on position bias, and seeking potential mitigation strategies.
@@
LLM Agreement Analysis: We complement our investigation of position bias with an agreement and disagreement analysis among LLM judges. Rather than focusing exclusively on overall consistency or fairness, we examine how LLM judges converge and diverge in their assessments of individual instances. Agreement analysis quantifies the percentage of instances where two LLM judges mutually agree on the outcome. Disagreement analysis counts the number of choices deviating from the mode for each instance among all judges. This further complies a “distribution of disagreement” across the dataset.
Our findings reveal that, despite exhibiting similar overall $PC$ and $PF$ scores, judges vary significantly in their judgments on individual instance. Disagreement analysis, in particular, highlights instances where consensus is either easily or difficultly achieved, reflecting the inherent complexity of the judgment task. For example, as shown in Fig. 3, more than half of the dataset can be considered relatively easy to judge, as over 80% of all 15 LLM judges agree with each other on these instances (disagreement $\leq$ 3). Conversely, fewer than 2% of instances represent the likely especially-hard-to-judge cases where a majority of LLM judges fail to reach consensus (disagreement $\geq$ 8).
Figure 3: Distribution of disagreement on MTBench. The y-axis indicates the proportion of the dataset where the level of disagreement among LLM judges does not exceed a specific threshold.
Based on our observations of answer quality gaps and LLM agreement/disagreement patterns, this study offers practical insights for designing evaluator benchmarks that account for the varying difficulty levels of judgment tasks. Specifically, the most challenging instances to evaluate are characterized by: (1) frequent disagreements among LLM judges, (2) closely matched win rates and minimal quality gaps among candidate models, and (3) significant position bias exhibited by the majority of LLM judges. Further discussions and analyses can be found in Appendix Sec. C.
@@ 6 Limitations
Despite proposing scalable metrics and investigating key factors influencing position bias, our study has several limitations.
First, we evaluated only 12 commercial closed-source LLM judges for pairwise settings and 4 for list-wise paradigms across two benchmarks, limiting the list-wise comparisons to three-candidate lists. Among open-source models, we assessed only three Llama 3.1+ models of varying sizes, as earlier versions or other smaller models lacked sufficient context window lengths, making them unsuitable given the length of evaluation instances in our study. Additionally, while we used original benchmark prompt templates, exploring alternative prompting techniques could offer further insights. Future work could expand on this by incorporating more models, tasks, prompting strategies, and larger list-wise candidate pools to enhance the generalizability of our findings.
Second, data accessibility limitations prevented a direct analysis of Judge-level factors like architecture and parameter size of closed-source models. Instead, we approximated these factors by grouping models by family properties. While open-source models offer accessible architectural details for deeper analysis, our assessment of only three Llama models may not provide sufficient evidence for broader conclusions. Additionally, our analyses were conducted post hoc, relying on completed judgments before analysis. Future work could explore methods to estimate or control these factors pre-judgment, reducing computational costs and enabling proactive mitigation strategies.
Lastly, our focus was on evaluating and understanding position bias rather than mitigating it. While our findings provide a foundation for effective mitigation, further research is needed to address issues like maintaining consistency and fairness when answer quality gaps are minimal, where position bias is most pronounced. Multivariate analyses exploring interactions between factors like prompt length, task complexity, and answer quality gaps could also yield deeper insights and enhance mitigation approaches.
@@ Appendix A Reproducibility
Our experiments were conducted primarily using API access, with a total cost of approximately 3,000 USD. The code repository for reproducing our results is available:https://anonymous.4open.science/r/Position-Bias-Analyzer-Demo-F7E3
@@ Appendix C LLM Agreement Analysis
Figure 4: Figures (a) to (c) are mutual agreement heatmap of LLM judges on MTBench and DevBench, where (b) is the agreement computation excluding the tie option {C}. Higher mutual agreement between two LLM judges is marked with brighter color. Figure (d), like Figure 3, is the distribution of disagreement on DevBench.
@@ C.1 Mutual Agreement & Familial Property
We compute the LLM judges’ mutual agreement on the instances to explore how “alike” or consistent they are across a set of evaluations. We denote two judges agree on an instance if their judgment choices are identical. Then the mutual agreement between two LLM judges on a benchmark is defined as the proportion of their agreed instances. Figs. 4(a) and (c) displays the mutual agreement heatmap for all judges on MTBench and DevBench, respectively. For MTBench that utilizes the 3-option mode, we also consider the “without tie" agreement since two judges are less disagreed when one chooses {C} while the other prefers a certain solution, compared to the case when they prefer different solutions. The “without tie” agreement heatmap of the twelve judges on MTBench is explored in Fig. 4(b).
@@
Llama models demonstrate a high agreement among capable family members (Llama-3.3-70B and Llama-405B) and with GPT series. However, significantly smaller and less capable model like Llama-3.1-8B does not strong agree with them.
@@ C.2 Disagreement & Benchmark Design Insight
Since the mutual agreement between LLM judges is not perfect and usually a considerable proportion of instances are difficult for them to reach a consensus, disagreement analysis becomes crucial and insightful. Therefore, we define the disagreement of an evaluation instance to be the number of judgments different from the majority. By this definition, an instance with all judges reaching a consensus on the better solution will have a disagreement of 0; in contrast, an instance where judgments are widely varied will result in a high disagreement. For our study where 15 judges are investigated, the maximum disagreement of an MTBench instance is 10, accounting for the 5{A}-5{B}-5{C} choice pattern by 3-option mode. On the other hand, for DevBench instances, since Gemini-1.5-flash is excluded due to insufficient data (as shown by high error rates in Tabel 2), the maximum possible disagreement for the remaining 14 judges is 7, representing the 7{A}-7{B} judgment distribution for the 2-option mode.
The distributions of instances with different disagreement values on MTBench and DevBench are shown in Fig. 3 and Fig. 4(d), respectively. From our disagreement analysis, at least 75% of the judges reached a choice consensus on more than half of the instances on both benchmarks. These are likely easy-to-evaluate instances, and the reliability of LLM judgments is enhanced by majority voting. In comparison, the instances with the highest disagreement are likely the ones that are difficult to evaluate and where the position bias is most likely to occur. However, luckily, these instances are rare, occupying only less than 5% for both benchmarks respectively. In other words, majority voting of multiple capable LLM judges could be practically useful for over 95% of evaluation instances on both benchmarks.
Moreover, if we roughly consider the disagreement value of instances as their difficulty for judging, then Fig. 3 and Fig. 4(d) exhibit a balanced distribution of instances with varied difficulty. This is because, except for the instances with the highest disagreement, the numbers of other instances with varied disagreement do not vary significantly, indicating a smoothly increasing difficulty curve across the benchmark datasets.
@@
Appendix D More Results of Position Bias and Answer Quality Gap Measurement
D.1 MTBench
@@ D.2 DevBench
Figure 6: Baseline comparisons of position bias of LLM judges across tasks on DevBench. An asterisk marks the statistical significance by Student’s t-tests. Figure (a) selects GPT-4 as the baseline, where asterisk demonstrates signficantly better or worse $PC$ of other models compared to it. Figure (b) utilizes an absolute $PF$ baseline of 0 and depicts preference fairness performances of LLM judges across tasks. Similar to findings on MTBench, position bias is significantly judge-dependent and task-dependent on DevBench as well.
Appendix E Variable Selection and Tests
E.1 Bidirectional Stepwise Regression with AIC
@@ E.2 Test results
Appendix F Experiment Settings
@@
F.1 Judges, Candidates, and Tasks